Jen butler has been pricing speed pass train fares for a group trip to new york. Three adults and four children must pay ree adults and four children must pay $89. Two adults and tree children must pay $63. Find the price of the adult ticket and the price of the childs ticket.
Answer by addingup(3677)

3a+4c = 89 multiply times 2
2a+3c = 63 multiply times 3
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Multiply, then subtract
6a+8c = 178
-
6a+9c = 189
------------
. -c = -11 divide both sides by -1
c = 11 This is the cost of a child's ticket. And the adult:
3a+4(11) = 89
3a+44 = 89
3a = 45
a = 15 is the cost of an adult ticket
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
3 adults*15 = 45
4 children*11= 44
Total. . . . 89 Correct
And:
2 adults*15 = 30
3 children*11= 33
Total. . . . . 63 Also correct